# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ Appendix A Detailed Proofs
where the last inequality follows from the connection condition in Algorithm 1. Therefore, $R_{\hat{\gamma}}(u)$ includes users that share the same preference vector as user $u$, while $W_{\hat{\gamma}}(u)$ includes users whose preference vectors differ from that of $u$ by at most $\hat{\gamma}=\gamma+\varepsilon$.
We now turn to bounding the suboptimality term $\text{SubOpt}(u,\pi,A_{\text{test}})=\left\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\right\rangle-\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$. For notational simplicity, we let:
$V=V_{\hat{\gamma}}(u),\quad R=R_{\hat{\gamma}}(u),\quad W=W_{\hat{\gamma}}(u),\quad W^{j}=W_{\hat{\gamma}}^{j}(u),\quad\tilde{J}=J_{\hat{\gamma}}(u).$
@@
Applying the Cauchy–Schwarz inequality, we obtain from Equation 10:
@@
$+\left\|\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\bm{\theta}^{j}-\bm{\theta}_{u})\right\|_{2}$ (11)
We now bound each term using matrix norm properties and known results from Equation 11:
$\left\langle\hat{\bm{\theta}}^{V}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle\leq\;$ $\frac{\beta(V,\delta)}{\sqrt{\lambda_{\min}(V)}}+\frac{\lambda|V|}{\lambda_{\min}(V)}+\frac{\lambda_{\max}(W)}{\lambda_{\min}(V)}\cdot\hat{\gamma}$
[…]
@@
Here, Equation 9 follows from the decomposition $M^{V}=M^{R}+\sum_{j\in\tilde{J}}M^{W^{j}}$; Equation 10 uses the fact that $M^{\mathcal{S}}=\lambda|\mathcal{S}|I+\sum_{v\in\mathcal{S}}\sum_{i=1}^{N_{v}}\bm{a}_{v}^{i}(\bm{a}_{v}^{i})^{\top}$ and the definition of reward noise $\varepsilon_{v}^{i}$. Inequality (11) is due to the Cauchy–Schwarz inequality. The first inequality in (12) uses eigenvalue bounds and Theorem 1 in [Abbasi-Yadkori et al., 2011], while the second inequality follows from bounding eigenvalues via the number of samples and applying Lemma B.2.
@@
$-\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle\leq\Psi-\left\langle\hat{\bm{\theta}}^{V},\bm{a}_{\text{test}}\right\rangle.$
@@
$\text{SubOpt}(u,\pi,A_{\text{test}})$ $=\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\rangle-\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle$
@@
$\leq 2\Psi+\frac{\beta(V,\delta)}{\sqrt{\lambda_{a}N^{V}}}$
with probability at least $1-2\delta$, which completes the proof of Lemma A.1. ∎
Proof of Lemma A.2.
From Lemma A.1 we know that
$\textnormal{SubOpt}(u_{\text{test}},\text{Off-NCLUB},A_{\text{test}})$ $\leq\tilde{O}\left(\frac{\sqrt{d}}{\sqrt{N^{{V}_{\hat{\gamma}}(u_{\text{test}})}}}+\eta_{{W}_{\hat{\gamma}}(u_{\text{test}})}\hat{\gamma}\right)$
$\leq\tilde{O}\left(\sqrt{\frac{d}{N^{R_{\hat{\gamma}}(u_{\text{test}})}}}+\eta_{{W}_{\hat{\gamma}}(u_{\text{test}})}\cdot\hat{\gamma}\right)$ (13)
which ends the proof of Lemma A.2.
Proof of Lemma A.3.
We now proceed to prove Lemma A.3.
Proof of $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}})$.
We begin by analyzing the estimation set $\mathcal{R}_{\hat{\gamma}}(u)$. To establish this subset relation, it suffices to show that, for
$\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{2\max\{2,d\}}},$
any user $v$ such that $\bm{\theta}_{u}=\bm{\theta}_{v}$, $K_{u}+K_{v}\leq\alpha_{r}\hat{\gamma}$ and $N_{v}\geq N_{\min}$ is included in $R_{\hat{\gamma}}(u)$ with high probability when $N_{u}\geq N_{\min}$.
Note that the condition $K_{u}+K_{v}<\frac{\hat{\gamma}}{2(\alpha+1)\sqrt{2\max\{2,d\}}}$ implies
$(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$ $\leq\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\lambda_{a}N_{u}}}+\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\lambda_{a}N_{v}}}$
$<\hat{\gamma},$
under the parameter selections for $\lambda$ and $\delta$ such that
Given that $\bm{\theta}_{u}=\bm{\theta}_{v}$, we further observe:
$\hat{\gamma}$ $>(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$
with probability at least $1-2\delta$, by Lemma A.1. This directly implies that the connection condition in Algorithm 1 is satisfied, and therefore user $v$ is included in $R_{\hat{\gamma}}(u)$.
Proof of $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}})$.
It suffices to show that if $K_{u}+K_{v}>\frac{\varepsilon}{\alpha-1}$ and $\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\geq\gamma$, then user $v$ is not included in $\mathcal{W}_{\hat{\gamma}}(u)$ with probability at least $1-2\delta$ when $\min\{N_{u},N_{v}\}\geq N_{\min}$.
Observe that the condition $K_{u}+K_{v}>\frac{\varepsilon}{\alpha-1}$ implies:
$(\alpha-1)(\text{CI}_{u}+\text{CI}_{v})$ $\geq(\alpha-1)\left(\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{u}}}+\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{v}}}\right)$
$>\varepsilon\geq\hat{\gamma}-\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
Therefore, we can deduce that:
$\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\|_{2}+\alpha(\text{CI}_{u}+\text{CI}_{v})$ $\geq\alpha(\text{CI}_{u}+\text{CI}_{v})+\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}-\text{CI}_{u}-\text{CI}_{v}$
$>\hat{\gamma},$
which means that the connection condition in Algorithm 1 does not hold. Hence, user $v$ will not be connected to $u$, and thus not included in $\mathcal{W}_{\hat{\gamma}}(u)$. Therefore, the superset condition holds with probability at least $1-2\delta$, provided $\alpha_{w}\in\left(0,\frac{1}{\alpha-1}\right)$.
Proof of $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}$.
Since we have shown that $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}})$ and $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}})$, the denominator in the definition of $\eta_{\mathcal{W}_{\hat{\gamma}}(u)}$ is larger while the numerator is at least as large, leading to:
$\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}.$
This concludes the proof of Lemma A.3.
We are now ready to prove Theorem 4.2. From Lemma A.2, we know that the suboptimality gap increases with the size of $W_{\hat{\gamma}}(u_{\text{test}})$ and decreases with the size of $R_{\hat{\gamma}}(u_{\text{test}})$. Therefore, by conservatively replacing $R_{\hat{\gamma}}(u_{\text{test}})$ with its subset estimate $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$, and replacing $W_{\hat{\gamma}}(u_{\text{test}})$ with its superset estimate $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, we can ensure a valid high-probability upper bound on the suboptimality gap.
A.2 Proof of Lemma 4.7
Proof of Lemma 4.7.
@@
We will show that any user included in $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, as defined in Theorem 4.2, also satisfies the condition in Lemma 4.7. For simplicity, we denote $u=u_{\text{test}}$ in the rest of the proof.
@@
Furthermore, using the definition of $K_{u}$ and $K_{v}$, we obtain:
$\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq 2(\alpha+1)\sqrt{\max\{2,d\}}(K_{u}+K_{v}).$
@@
$K_{u}+K_{v}\geq\frac{\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}}{2(\alpha+1)\sqrt{\max\{2,d\}}}.$
Thus, by letting $\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{\max\{2,d\}}}$, we obtain the desired result:
$K_{u}+K_{v}\geq\alpha_{r}\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
A.3 Proof of Lemma 4.8
Proof of Lemma 4.8.
The proof of Lemma 4.8 follows from the fact that for all $v\in M(u_{\textnormal{test}})$, we have
@@
$\mathcal{F}_{u}(\delta)$ =$=\left\{\left\|\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right\|_{\tilde{M}_{u}}\leq\beta(\tilde{N}_{u},\delta)\right\},\quad\forall u\in\mathcal{V},$
$\mathcal{F}^{\prime}_{u}(\delta)$ =$=\left\{\left\|\bm{\theta}_{u}-\hat{\bm{\theta}}_{u}\right\|_{M_{u}}\leq\beta(N_{u},\delta)\right\},\quad\forall u\in\mathcal{V}.$ (16)
We first verify that event $\mathcal{E}$ holds with high probability. According to Lemma B.1, the event $\mathcal{F}^{\prime}_{u}(\delta/(2U))$ holds for all $u\in\mathcal{V}$ with probability at least $1-\frac{1}{2}\delta$. Under this event, we have:
@@
$\leq\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\lambda_{a}N_{u}/2}}$
@@
$N_{u}\geq\max\left\{\frac{512d}{\gamma^{2}\lambda_{a}}\log\left(\frac{2U}{\delta}\right),\ \frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)\right\}.$
The last inequality in (17) follows from Lemma B.4.
Now consider the connection rule in Algorithm 2, specifically the separation condition in Equation 7. For any users $u_{1},u_{2}\in\mathcal{V}$, the algorithm distinguishes them into different clusters if
@@
$\text{SubOpt}(u,\pi,A_{\text{test}})$ $\leq\left(\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right)^{\top}\bm{a}_{\text{test}}^{*}+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}$
@@
From Lemma B.2, we know
$\lambda_{\min}(\tilde{M}_{u})\geq\lambda+\frac{\lambda_{a}N^{\mathcal{V}(j_{u})}}{2},$
@@
$\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}\leq\frac{1}{\sqrt{\lambda+\lambda_{a}N^{\mathcal{V}(j_{u})}/2}}.$
@@
$\text{SubOpt}(u_{\text{test}},\text{Off-CLUB},A_{\text{test}})\leq\frac{\sqrt{d\log\left(1+\frac{N^{\mathcal{V}(j_{u_{\text{test}}})}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\lambda_{a}N^{\mathcal{V}(j_{u_{\text{test}}})}/2}},$ (19)
which holds with probability at least $1-2\delta$. Using the order $\tilde{O}$ to denote Equation 19 ends our proof. ∎
A.5 Proof of Theorem 6.1
Proof of Theorem 6.1.
@@
$N^{\mathcal{V}(j_{u_{\text{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\text{test}}})}N_{v}.$
@@
Appendix B Technical Lemmas

### v2

@@ Appendix C Detailed Proofs
where the last inequality follows from the connection condition in Algorithm 1. Therefore, $\mathcal{R}_{\hat{\gamma}}(u)$ includes users that share the same preference vector as user $u$, while $\mathcal{W}_{\hat{\gamma}}(u)$ includes users whose preference vectors differ from that of $u$ by at most $\hat{\gamma}=\gamma+\varepsilon$.
Next, we prove Lemma˜4.2, beginning with the representation of $\mathcal{R}_{\hat{\gamma}}(u)$ and then proceeding to that of $\mathcal{W}_{\hat{\gamma}}(u)$. Throughout, we focus on the case where $N_{u}\geq N_{\min}$.
Proof of $\mathcal{R}_{\hat{\gamma}}(u)$.
We begin by analyzing the set $\mathcal{R}_{\hat{\gamma}}(u)$. Specifically, it suffices to show: With probability at least $1-2\delta$, (i) for $\alpha_{r1}=\tfrac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}}$, any user $v$ with $\bm{\theta}_{u}=\bm{\theta}_{v}$, $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}\leq\alpha_{r1}\hat{\gamma}$, and $N_{v}\geq N_{\min}$ is included in $\mathcal{R}_{\hat{\gamma}}(u)$ when $N_{u}\geq N_{\min}$; (ii) for $\alpha_{r2}=\tfrac{\sqrt{\tilde{\lambda}_{a}}}{2\alpha\sqrt{\log(2U/\delta)}}$, any user $v$ with $\bm{\theta}_{u}\neq\bm{\theta}_{v}$, or $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}>\alpha_{r2}\hat{\gamma}$, or $N_{v}<N_{\min}$ cannot be included in $\mathcal{R}_{\hat{\gamma}}(u)$.
Proof of (i). Note that the condition $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}\leq\alpha_{r1}\hat{\gamma}$ implies
$(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$ $\leq\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\tilde{\lambda}_{a}N_{u}}}+\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\tilde{\lambda}_{a}N_{v}}}$
$<\hat{\gamma}$,
under the parameter selections for $\lambda$ and $\delta$ such that
Given that $\bm{\theta}_{u}=\bm{\theta}_{v}$, we further observe:
$\hat{\gamma}$ $>(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$
with probability at least $1-2\delta$. This directly implies that the connection condition in Algorithm 1 is satisfied, and therefore user $v$ is included in $\mathcal{R}_{\hat{\gamma}}(u)$.
Proof of (ii). If user $v$ satisfies either $\bm{\theta}_{u}\neq\bm{\theta}_{v}$ or $N_{v}<N_{\min}$, it is straightforward that $v\notin\mathcal{R}_{\hat{\gamma}}(u)$. We therefore focus on showing that the condition $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}>\alpha_{r2}\hat{\gamma}$ also implies $v\notin\mathcal{R}_{\hat{\gamma}}(u)$.
Note that the condition that $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}>\alpha_{r2}\hat{\gamma}$ implies
$\alpha(\text{CI}_{u}+\text{CI}_{v})>2\alpha\left(\sqrt{\frac{\log\left(\frac{2U}{\delta}\right)}{\tilde{\lambda}_{a}N_{u}}}+\sqrt{\frac{\log\left(\frac{2U}{\delta}\right)}{\tilde{\lambda}_{a}N_{v}}}\right)>\hat{\gamma}$.
Hence the connecting condition in Equation˜2 does not hold due to the positive value of $\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}$ and $v\notin\mathcal{R}_{\hat{\gamma}}(u)$ holds.
Proof of $\mathcal{W}_{\hat{\gamma}}(u)$.
First it is trivial that $\alpha_{w}>0$ holds. As we have proved that for any $v\in\mathcal{W}_{\hat{\gamma}}(u)$ it holds that $\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}$, therefore it suffices to show that if $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}>\frac{\varepsilon\sqrt{\tilde{\lambda}_{a}}}{2(\alpha-1)\sqrt{\log(2U/\delta)}}$ and $\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\geq\gamma$, then user $v$ is not included in $\mathcal{W}_{\hat{\gamma}}(u)$ with probability at least $1-2\delta$ when $\min\{N_{u},N_{v}\}\geq N_{\min}$.
Observe that the condition $1/\sqrt{N_{u}}+1/\sqrt{N_{v}}>\frac{\varepsilon\sqrt{\tilde{\lambda}_{a}}}{2(\alpha-1)\sqrt{\log(2U/\delta)}}$ implies:
$(\alpha-1)(\text{CI}_{u}+\text{CI}_{v})$ $\geq(\alpha-1)\left(\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\tilde{\lambda}_{a}N_{u}}}+\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\tilde{\lambda}_{a}N_{v}}}\right)$
$>\varepsilon\geq\hat{\gamma}-\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
Therefore, we can deduce that:
$\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}+\alpha(\text{CI}_{u}+\text{CI}_{v})$ $\geq\alpha(\text{CI}_{u}+\text{CI}_{v})+\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}-\text{CI}_{u}-\text{CI}_{v}$
$>\hat{\gamma}$,
which means that the connection condition in Equation˜2 does not hold. Consequently, user $v$ is not connected to $u$, and therefore not included in $\mathcal{W}_{\hat{\gamma}}(u)$ with probability at least $1-2\delta$, provided that $\alpha_{w}\in\left(0,\tfrac{\sqrt{\tilde{\lambda}_{a}}}{2(\alpha-1)\sqrt{\log(2U/\delta)}}\right)$.
C.2 Proof of Theorem˜4.4
Proof of Theorem˜4.4.
To prove Theorem˜4.4, we first recall and several key notations and introduce some new notations in Table 5.
Table 5: Notations in Proofs
Notation Description
$\mathcal{V}_{\hat{\gamma}}(u)$ Set consisting of user $u$ and its neighbors in the graph $\mathcal{G}_{\hat{\gamma}}$.
$\mathcal{R}_{\hat{\gamma}}(u)$ Set of users in $\mathcal{V}_{\hat{\gamma}}(u)$ that share the same preference vector as user $u$.
$\mathcal{W}_{\hat{\gamma}}(u)$ Set of users in $\mathcal{V}_{\hat{\gamma}}(u)$ that do not share the same preference vector as user $u$.
$\mathcal{W}_{\hat{\gamma}}^{j}(u)$ Set of users in $\mathcal{W}_{\hat{\gamma}}(u)$ who belong to cluster $\mathcal{V}(j)$.
$\mathcal{J}_{\hat{\gamma}}(u)$ Set of indices of distinct clusters represented in $\mathcal{W}_{\hat{\gamma}}(u)$.
To characterize the contributions of different types of neighbors in the estimation process, we define the weighted fraction of samples from $\mathcal{R}_{\hat{\gamma}}(u)$ relative to $\mathcal{V}_{\hat{\gamma}}(u)$ as:
$\eta_{\mathcal{R}_{\hat{\gamma}}(u)}=\frac{|\mathcal{R}_{\hat{\gamma}}(u)|\lambda+N_{\mathcal{R}_{\hat{\gamma}}(u)}}{|\mathcal{V}_{\hat{\gamma}}(u)|\lambda+N_{\mathcal{V}_{\hat{\gamma}}(u)}}.$
Here, $\lambda$ is the regularization parameter used in constructing each user’s matrix $M_{u}$. Analogous definitions apply to $\eta_{\mathcal{W}_{\hat{\gamma}}(u)}$ and $\eta_{\mathcal{W}_{\hat{\gamma}}^{j}(u)}$, and by construction, we have:
$\eta_{\mathcal{R}_{\hat{\gamma}}(u)}+\eta_{\mathcal{W}_{\hat{\gamma}}(u)}=1,$
since $\mathcal{V}_{\hat{\gamma}}(u)=\mathcal{R}_{\hat{\gamma}}(u)\cup\mathcal{W}_{\hat{\gamma}}(u)$ and the two sets are disjoint.
We now turn to bounding the suboptimality term $\text{SubOpt}(u,\pi,\mathcal{A}_{\text{test}})=\left\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\right\rangle-\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$. Firstly, the case where $N_{u}<N_{\min}$ is trivial since the problem can be simplified to an offline linear contextual bandit problem in [Li et al., 2022] as the algorithm only makes use of the data from user $u$ itself. For the non-trivial case where $N_{u}\geq N_{\min}$, recall that in Section˜C.1 we proved that $\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{2}\leq\text{CI}_{v}$ for any user $v$ with probability at least $1-\frac{3}{2}\delta$, and $\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}\leq\hat{\gamma}$ holds for any connected users $(u_{1},u_{2})$.
For notational simplicity, we let:
$V=\mathcal{V}_{\hat{\gamma}}(u),\quad R=\mathcal{R}_{\hat{\gamma}}(u),\quad W=\mathcal{W}_{\hat{\gamma}}(u),\quad W^{j}=\mathcal{W}_{\hat{\gamma}}^{j}(u),\quad\tilde{J}=\mathcal{J}_{\hat{\gamma}}(u).$
@@
Applying the Cauchy–Schwarz inequality, we obtain from Equation˜9:
@@
$+\left\|\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\bm{\theta}^{j}-\bm{\theta}_{u})\right\|_{2}\cdot\left\|\bm{a}_{\text{test}}\right\|_{2}$ (10)
We now bound each term using matrix norm properties and known results from Equation˜10:
$\left\langle\hat{\bm{\theta}}^{V}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle\leq\;$ $\frac{\beta(V,\delta)}{\sqrt{\lambda_{\min}(V)}}+\frac{\lambda|V|}{\lambda_{\min}(V)}+\frac{\lambda_{\max}(W)}{\lambda_{\min}(V)}\cdot\hat{\gamma}$ (11)
$\leq\;$ $\frac{\sqrt{2}\beta(V,\delta)}{\sqrt{\tilde{\lambda}_{a}N_{V}}}+\frac{2\lambda U}{\tilde{\lambda}_{a}N_{V}}+\frac{2\eta_{W}}{\tilde{\lambda}_{a}}\cdot\hat{\gamma}\eqqcolon\Psi$ (12)
@@
Here, Equation˜8 follows from the decomposition $M^{V}=M^{R}+\sum_{j\in\tilde{J}}M^{W^{j}}$; Equation˜9 uses the fact that $M^{\mathcal{S}}=\lambda|\mathcal{S}|I+\sum_{v\in\mathcal{S}}\sum_{i=1}^{N_{v}}\bm{a}_{v}^{i}(\bm{a}_{v}^{i})^{\top}$ and the definition of reward noise $\varepsilon_{v}^{i}$. Equation˜10 is due to the Cauchy–Schwarz inequality. The first inequality in Equation˜12 uses eigenvalue bounds and Theorem 1 in [Abbasi-Yadkori et al., 2011], while the second inequality follows from bounding eigenvalues via the number of samples and applying ˜3.1, Lemma J.1 in Wang et al. [2023a] and Lemma 7 in Li and Zhang [2018].
@@
$\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle\leq\Psi-\left\langle\hat{\bm{\theta}}^{V},\bm{a}_{\text{test}}\right\rangle.$
@@
$\text{SubOpt}(u,\pi,\mathcal{A}_{\text{test}})$ $=\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\rangle-\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle$
@@
$\leq 2\Psi+\frac{\beta(V,\delta)}{\sqrt{\tilde{\lambda}_{a}N_{V}}}$
with probability at least $1-2\delta$. By omitting the logarithmic and constant terms in the above inequality we complete the proof of Theorem˜4.4. ∎
C.3 Proof of Lemma B.2
Proof of Lemma B.2.
@@
We will show that any user included in $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, as represented in Lemma˜4.2, also satisfies the condition in Lemma B.2. For simplicity, we denote $u=u_{\text{test}}$ in the rest of the proof.
@@
Furthermore, using the definition of $\text{CI}_{u}$ and $\text{CI}_{v}$, we obtain:
$\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\frac{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}}{\sqrt{\tilde{\lambda}_{a}}}\left(\frac{1}{\sqrt{N_{u}}}+\frac{1}{\sqrt{N_{v}}}\right).$
@@
$\frac{1}{\sqrt{N_{u}}}+\frac{1}{\sqrt{N_{v}}}\geq\frac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\log(2U/\delta)\max\{2,d\}}}\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
Thus, by letting $\alpha_{w}^{\prime}=\frac{\sqrt{\tilde{\lambda}_{a}}}{4(\alpha+1)\sqrt{\max\{2,d\}\log(2U/\delta)}}$, we obtain the desired result:
$\frac{1}{\sqrt{N_{u}}}+\frac{1}{\sqrt{N_{v}}}\geq\alpha_{w}^{\prime}\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
Then for the second case that $\Gamma(u_{\text{test}},v)\geq\hat{\gamma}$, it can be proved that
$\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}\geq\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}-\alpha\left(\text{CI}_{u}+\text{CI}_{v}\right)\geq\hat{\gamma}$.
Thus $v\notin\mathcal{W}_{\hat{\gamma}}(u)$ holds. This ends the proof of Theorem˜B.2. ∎
C.4 Proof of Lemma B.4
Proof of Lemma B.4.
The proof of Lemma B.4 follows from the fact that for all $v\in M(u_{\textnormal{test}})$, as $M(u_{\text{test}})$ denotes the users belong to different clusters of $u_{\text{test}}$. By combining Equation˜4, we have
@@
$\mathcal{F}_{u}(\delta)$ $=\left\{\left\|\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right\|_{\tilde{M}_{u}}\leq\beta(\tilde{N}_{u},\delta)\right\},\quad\forall u\in\mathcal{V},$
$\mathcal{F}^{\prime}_{u}(\delta)$ $=\left\{\left\|\bm{\theta}_{u}-\hat{\bm{\theta}}_{u}\right\|_{M_{u}}\leq\beta(N_{u},\delta)\right\},\quad\forall u\in\mathcal{V}.$ (15)
We first verify that event $\mathcal{E}$ holds with high probability. According to Lemma D.1, the event $\mathcal{F}^{\prime}_{u}(\delta/(2U))$ holds for all $u\in\mathcal{V}$ with probability at least $1-\frac{1}{2}\delta$. Under this event, we have:
@@
$\leq\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\tilde{\lambda}_{a}N_{u}/2}}$
@@
$N_{u}\geq\max\left\{\frac{512d}{\gamma^{2}\tilde{\lambda}_{a}}\log\left(\frac{2U}{\delta}\right),\ \frac{16}{\tilde{\lambda}_{a}^{2}}\log\left(\frac{8dU}{\tilde{\lambda}_{a}^{2}\delta}\right)\right\}.$
The last inequality in Equation˜16 follows from Lemma D.5.
Now consider the connection rule in Algorithm 2, specifically the separation condition in Equation˜6. For any users $u_{1},u_{2}\in\mathcal{V}$, the algorithm distinguishes them into different clusters if
@@
$\text{SubOpt}(u,\pi,\mathcal{A}_{\text{test}})$ $\leq\left(\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right)^{\top}\bm{a}_{\text{test}}^{*}+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}$
@@
From ˜3.1, Lemma J.1 in Wang et al. [2023a] and Lemma 7 in Li and Zhang [2018], we know
$\lambda_{\min}(\tilde{M}_{u})\geq\lambda+\frac{\tilde{\lambda}_{a}N_{\mathcal{V}(j_{u})}}{2},$
@@
$\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}\leq\frac{1}{\sqrt{\lambda+\tilde{\lambda}_{a}N_{\mathcal{V}(j_{u})}/2}}.$
@@
$\text{SubOpt}(u_{\text{test}},\text{Off-CLUB},\mathcal{A}_{\text{test}})\leq\frac{\sqrt{d\log\left(1+\frac{N_{\mathcal{V}(j_{u_{\text{test}}})}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\tilde{\lambda}_{a}N_{\mathcal{V}(j_{u_{\text{test}}})}/2}},$ (18)
which holds with probability at least $1-2\delta$. Using the order $\tilde{O}$ to denote Equation˜18 ends our proof. ∎
C.6 Proof of Theorem 5.4
Proof of Theorem˜5.4.
@@
$N_{\mathcal{V}(j_{u_{\text{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\text{test}}})}N_{v}.$
@@
Appendix D Technical Lemmas